# Privacy Bills of Materials: A Transparent Privacy Information Inventory for Collaborative Privacy Notice Generation in Mobile App Development

### v1

@@ 5.1 User Study Design
5.2 Participant Recruitment
Following the participant recruitment processes in prior works in Usable Privacy and Security field [80, 54, 62], we published the survey via Qualtrics [16], and recruited participants through Prolific [15], a commonly used recruitment platform of sustainably gathering survey responses at scale. We use the participant screener in Prolific to recruit participants that meet our requirement discussed in Section 5.1.
We consider responses that are uncompleted or with a completion time of less than 2 minutes to be invalid. In total, we received 150 valid survey responses from participants with diverse backgrounds. The median time to complete the survey is 14 minutes and 33 seconds. Our participants span across 35 countries include 94 males and 55 females, with one participant preferring not to say, which conforms the reality that there are more male practitioners in software industry. We also include 3 demographic questions in the survey questionnaire to understand the roles of participants in software development, years of experience, and the size of the team they work in. An overview of the demographic information is presented in Table 3. The role with the highest proportion is junior developer (39%), while other participants reported to be UI designers (15%), project managers (14%), senior developers (13%), legal team members (9%), and other roles such as CEO. The team size among participants is also diverse. The most common team size is less than 10 people (59%), while there are also 9% of participants working in a team with more than 50 people. We believe our participants’ backgrounds are sufficiently diverse for our survey.
5.3 Result Analysis
The results of the survey are detailed in Table 4. We calculate the average agreement score and present the distribution of Likert-scale responses. Overall, participants largely agree with our statements regarding PriBOM. Below, we discuss the results in more detail.
5.3.1 Design of PriBOM
@@
5.3.2 Viewpoint Comparison Across Roles
@@
Similarly, UI designers scored an average of 3.82, while legal teams scored significantly higher at 4.31 on the statement concerning the impact of PriBOM on enhancing privacy awareness among team members (Figure 5(d)). This means that though legal experts may see PriBOM as a critical tool for privacy practicing, UI designers might appreciate its role but to a lesser extent, as their direct interaction with privacy as a compliance or educational tool is less pronounced. Distinct differences in perception between the project managers and UI Designers were observed in the disclosure-behaviour alignment aspect of PriBOM (Figure 5(c)). The project managers, who need to ensure the development aligns with business and compliance objectives, show a higher agreement score of 4.19, suggesting a strong recognition of the utility of PriBOM in streamlining and clarifying the alignment between privacy disclosures and the actual behaviors. Regarding the maintenance of a record of privacy-related updates (Figure 5(b)), UI designers gave a lower score of 3.64, compared to senior developers who scored it at 4.16. Most discussions of privacy on Reddit were triggered by external events  [50], including updates of third-party code that is used. While senior developers are likely to value features that aid in documenting and archiving changes, which supports better version control and compliance tracking, UI designers may not perceive it as impactful to their role.
@@
Figure 5: Examples of the differences of agreement scores regarding survey statements among different roles.
@@
5.3.3 Perspectives Held by Non-technical Roles
@@
6 Conclusion
Privacy regulations commonly require developers to provide authentic and comprehensive privacy notices, e.g., privacy policies or labels, in communicating their apps’ privacy practices. However, developers often struggle to create and maintain accurate privacy notices, especially for sophisticated apps with complex features and large development teams. We introduce PriBOM (Privacy Bills of Materials) as a systematic and collaborative solution for the development team to craft accurate privacy notices. As a privacy information inventory, we leverage static analysis and privacy notice analysis techniques to implement PriBOM. The results of our human evaluation underscore the usefulness and practicability of PriBOM in fostering a privacy-conscious development workflow. Lastly, we discuss in depth the implications of our results, including differences in views held by different roles regarding PriBOM and privacy notice generation, especially non-technical roles.

### v2

@@ 5.3. Ethics Considerations
5.4. Result Analysis
The results of the survey are detailed in Table 4. We calculate the average agreement score and present the distribution of Likert-scale responses. Overall, participants largely agree with our statements regarding PriBOM. For open-ended responses, we employ the methodology of deductive thematic analysis (Braun and Clarke, 2006; Crabtree and Miller, 1992) to identify key insights. Given that the free-text questions are already structured around specific sections of our survey (e.g., $\textit{Q}_{1}$ corresponds to Design of the Codebase and Permission Section in Table 4), we align our qualitative analysis with predefined two-layer (sections and statements) a priori rather than emerging new codes and themes from responses. Below, we discuss the results in more detail.
5.4.1. Design of PriBOM
@@
5.4.2. Viewpoint Comparison Across Roles
@@
Similarly, UI designers scored an average of 3.82, while legal teams scored significantly higher at 4.31 on the statement concerning the impact of PriBOM on enhancing privacy awareness among team members (Figure 5(d)). This means that though legal experts may see PriBOM as a critical tool for privacy practicing, UI designers might appreciate its role but to a lesser extent, as their direct interaction with privacy as a compliance or educational tool is less pronounced. Distinct differences in perception between the project managers and UI Designers were observed in the disclosure-behaviour alignment aspect of PriBOM (Figure 5(c)). The project managers, who need to ensure the development aligns with business and compliance objectives, show a higher agreement score of 4.19, suggesting a strong recognition of the utility of PriBOM in streamlining and clarifying the alignment between privacy disclosures and the actual behaviors. Regarding the maintenance of a record of privacy-related updates (Figure 5(b)), UI designers gave a lower score of 3.64, compared to senior developers who scored it at 4.16. Most discussions of privacy on Reddit were triggered by external events  (Li et al., 2021a), including updates of third-party code that is used. While senior developers are likely to value features that aid in documenting and archiving changes, which supports better version control and compliance tracking, UI designers may not perceive it as impactful to their role.
@@
Figure 5. Examples of the differences of agreement scores regarding survey statements among different roles.
@@
5.4.3. Perspectives Held by Non-technical Roles
@@
6. Limitations
While we propose a pre-fill of PriBOM and access the usefulness, there are still limitations. First, the pre-fill generation is highly dependent on the performance of static analysis tools, and limited performance could result in restricted quality of pre-fill. Second, our user study design can be improved by involving a in-depth interview to mitigate the possibility of social desirability or acquiescence biases. Although we propose this promising privacy enhancing mechanism in software development process, real practicability needs to be further verified. Still, as the first of such kind, we believe this work can ignite deeper thinking on the aspect of software engineering at the systematic level, considering all practitioners to solve the long-standing privacy policy challenges.
7. Conclusion
Privacy regulations commonly require developers to provide authentic and comprehensive privacy notices, e.g., privacy policies or labels, in communicating their apps’ privacy practices. However, developers often struggle to create and maintain accurate privacy notices, especially for sophisticated apps with complex features and large development teams. We introduce PriBOM (Privacy Bills of Materials) as a systematic and collaborative solution for the development team to craft accurate privacy notices. As a privacy information inventory, we leverage static analysis and privacy notice analysis techniques to implement PriBOM. The role of PriBOM in enhancing privacy-related communication is well received with 83.33% agreement, underscoring the usefulness and practicability of PriBOM in fostering a privacy-conscious development workflow. Lastly, we discuss in depth the implications of our results, including differences in views held by different roles regarding PriBOM and privacy notice generation, especially non-technical roles.